Questions 1-5 Gigaland’s 2019 National Accounts

1. Personal taxes

320

2. Corporateincometaxes

70

3. Indirect taxes less subsidies

100

4. Undistributedcorporateprofits

120

5. Netincomeoffarmandunincorporatedbusiness

240

6. Wages, salaries, and supplementary labour income

1,000

7. Personal consumption expenditures

1,040

8. Capital consumption allowance (depreciation)

180

9. Exports

700

10. Profits of corporations and government enterprises before taxes

240

11.Interest and investment income

80

12. Dividends

90

13. Imports

640

14.Gross investment

340

15.Net investments from non-residents

160

16.Transfer payments to persons

140

17.Government purchases

440

  1. Gross Domestic Product (GDP) was
    a) $1,630; b) $1,680; c) $1,790; d) $1,880

  2. National Income was
    a) $1,550; b) $1,600; c) $1,750; d) $1,890

  3. Personal income was
    a) $1,320; b) $1,470; c) $1,500; d) $1,550

  4. Personal disposable income was
    a) $1,150; b) $1,230; c) $1,290; d) $1,380

  5. The net addition to the country’s stock of capital good was a) $160; b) $180; c) $210; d) $230
Answer #1

Answer:

1]

GDP = Consumption expenditures + government Purchase + net exports + investment.

= 1040 + 440 + (700 – 640) + 340

= $1880

Correct option: D] $1880

2]

NDP=GDP-depreciation = 1880 – 180 = 1700

NI=NDP + Corporate income taxes = 1700 – 70 = 1630

Correct option: A] $1630

3]

NNI = NDP - indirect taxes less subsidy - corporate tax

= 1700 - 100 - 70

= 1530

PI = NNI - Undistributed Corporate Profits + Transfer payments to persons

= 1530 - 120 + 140

= 1550

Correct option: A] $1550

4]

Personal disposable income = PI - personal income tax

= 1550 - 320

= 1230

Correct option: B] $1230

5]

Net investment = gross investment - depreciation

Net investment = 340 - 180 = $160 .

Therefore, the answer is A] $160.

The net addition to the country’s stock of capital good was $160